OpenStudy (anonymous):

Ben bought a digital camera for $90. The tax rate is 6%. Which proportion can be used to find the amount of tax Ben will pay?

OpenStudy (victoriasushchik):

Ben bought a digital camera for $90. The tax rate is 6%. Which proportion can be used to find the amount of tax Ben will pay? 90 + 90(0.06) = $95.4 (answer)

OpenStudy (victoriasushchik):

did that help?

OpenStudy (victoriasushchik):

x/90=6/100